For a Bose-Einstein system, the expression for the thermodynamic probability is:
W=π{(gi+Ni–1)!/(Ni !(gi–1)!)}
Derive an expression for the Bose-Einstein distribution function.
1
Expert's answer
2020-02-28T10:09:17-0500

As per the Bose-Einstein system,

The particles are indistinguishable.

Let total number of particle in the system is "n_i" ​ , energy "E_i"

​ and let the statistical weight factor(degeneracy) is "g_i"

​ .Now, the way of the distribution of "n_i" particle into "g_i" space,

So, the distribution "\dfrac{(n_i-g_i-1)!}{n_i(g_i-1)!}"

for "n_o" particle,

we know that thermodynamic probability "w=\dfrac{(n_o-g_o-1)!}{n_o(g_o-1)!}.\dfrac{(n_1-g_1-1)!}{n_1(g_1-1)!}.........\dfrac{(n_i-g_i-1)!}{n_i(g_i-1)!}"

So, we can write it as "w=\Pi\dfrac{(n_i-g_i-1)!}{n_i(g_i-1)!}"

Hence "\ln w=\sum {\ln((n_i-g_i-1)!)- \ln n_i-\ln(g_i-1)!}"